Factor the below mentioned question
a(x+y)-(x+y)^2
Can you please explain it by step by step
Thank you

Answer by ikleyn(52781)

.

In your post,  THERE  IS  NO  "below mentioned question".

What is below, is called "an expression" in Math . . .

-------------

a(x+y) - (x+y)^2 = (x+y)*(a-(x-y)) = (x+y)*(a-x-y).    ANSWER